# ToolACE-R: Model-aware Iterative Training and Adaptive Refinement for Tool Learning

### v2

@@ Abstract
Tool learning, which allows Large Language Models (LLMs) to leverage external tools for solving complex user tasks, has emerged as a promising avenue for extending model capabilities. However, existing approaches primarily focus on data synthesis for fine-tuning LLMs to invoke tools effectively, largely ignoring how to fully stimulate the potential of the model. In this paper, we propose ToolACE-R, a novel framework that includes both model-aware iterative training and adaptive refinement for tool learning. ToolACE-R features a model-aware iterative training procedure that progressively adjust training samples based on the model’s evolving capabilities to maximize its potential. Additionally, it incorporates self-refinement training corpus which emphasizes LLM’s ability to iteratively refine their tool calls, optimizing performance without requiring external feedback. Furthermore, we introduce adaptive self-refinement mechanism for efficient test-time scaling, where the trained model can autonomously determine when to stop the process based on iterative self-refinement. We conduct extensive experiments across several benchmark datasets, showing that ToolACE-R achieves competitive performance compared to advanced API-based models. The performance of tool invocation can be further improved efficiently through adaptive self-refinement. These results highlight the effectiveness and generalizability of ToolACE-R, offering a promising direction for more efficient and scalable tool learning.
1 Introduction
@@
Tool learning, which enables Large Language Models (LLMs) to leverage external tools to address complex user requirements, has gained increasing attention. With tool integration, LLMs can access up-to-date information, perform intricate computations, and utilize third-party services, significantly expanding their capabilities beyond simple natural language communication with humans [27]. While tool invocation requires the LLMs to demonstrate strong understanding, reasoning, and instruction-following skills, customized fine-tuning is currently the dominant approach for enabling models to call external tools [19, 20, 23, 25].
@@
Due to the limited availability of high-quality data, existing research has primarily focused on developing effective and efficient methods for data synthesis using advanced models [1, 19, 20, 35]. However, data synthesized by advanced models from different sources can lead to compatibility issues. Specifically, when synthesized samples exceed the model’s current knowledge, they may undermine the model’s performance or lead to hallucinations [12, 29]. As a result, determining which training data samples are appropriate for a given model remains an unresolved challenge.
@@
Additionally, the potential for maximizing a model’s intrinsic capabilities remains underexplored in the field of tool learning. On one hand, prior work has seldom investigated how to fully exploit existing training data through techniques such as data augmentation. On the other hand, scaling test-time computation – despite its demonstrated effectiveness in enhancing LLM reasoning [4, 32] – has received limited attention in tool learning scenarios. A promising approach involves iteratively refining the model’s outputs and selecting the majority answer [32]. However, user queries vary widely in complexity, ranging from simple to highly intricate. Scaling test-time computation indiscriminately is inefficient, as models can often correctly answer straightforward questions without the need for additional refinement. Therefore, an adaptive strategy for scaling test-time computation is needed to dynamically determine the appropriate level of computational effort required for each query.
@@
We have conducted extensive experiments on several representative tool-calling benchmarks, such as the Berkeley Function Call Leaderboard (BFCL) [38] and API-Bank [15], to assess the effectiveness and efficiency of ToolACE-R. Experimental results demonstrate that ToolACE-R achieves competitive performance compared to advanced API-based models such as GPT-4o, and can be further enhanced through adaptive self-refinement. Additional analyses highlight the contributions of the proposed modules and examine the generalizability of our method across different model backbones and sizes.
The key contributions of this work are summarized as follows:
@@ 2 Related Work
Tool learning methods generally fall into two categories, tuning-free and tuning-based approaches. Tuning-free methods employ prompting techniques without additional training [22, 8, 30]. A notable example method is ReAct [40], which can alternate between reasoning and action for complex tasks. Recent work [26] improves tool understanding by iteratively refining documentation. Tuning-based methods, which directly enhance tool calling ability via specialized fine-tuning, have recently gained much more attention [25, 31, 23, 33, 1]. A large amount of them focus on data synthesis for improved generation procedures [19, 20, 35], often overlooking the model’s inherent potential. While some others integrate tool feedback for refinement [7, 34], our approach is more streamlined, directly refining outputs in an iterative manner without external or any textual feedback to maximize model efficiency.
@@
Research indicates that selecting high-quality training samples is essential for fine-tuning LLMs [3]. A small set of high-quality data can effectively harness the model’s potential, rather than relying on large quantities of data [42, 18]. While earlier works primarily emphasize general data quality aspects, such as diversity and complexity, recent studies advocate for model-specific data selection [6, 14]. This kind of approach is based on the observation that data distributions that significantly deviate from the base model’s are challenging for the model to learn from and may even degrade the performance [29]. Our method aligns with this perspective, defining a metric with passrate@k to identify and select suitable training samples based on the model to be trained.
@@
Previous research has demonstrated that LLMs can refine their own outputs through either self-feedback [21, 36] or external feedback [28, 37]. Nevertheless, it remains challenging for LLMs to assess the correctness of their refined output autonomously [10], which means that scaling test-time compute with iterative refinement still often depend on post-processing techniques, such as self-consistency [32]. In this work, we propose a simple yet effective method that enables models to adaptively self-refine their outputs, thereby improving scaling efficiency.
@@
3 Methodology
3.1 Problem Formulation
Tool learning aims to address the following problem: given a user query $q$ and a set of candidate tools $T=\{t_{1},t_{2},\dots,t_{n}\}$, the objective is to generate the correct tool invocations. This includes selecting the most appropriate tools and extracting the suitable parameter values. Specifically, the goal is to produce:
@@ 3.1 Problem Formulation
Our ToolACE-R is trained with a model-aware iterative training pipeline and can iteratively self-refine during inference in an adaptive manner. Figure 1 shows the overall procedure.
3.2 Model-Aware Iterative Training
@@
To guide tool learning with finer granularity, we introduce a model-aware difficulty metric that quantifies the inherent challenge a specific training sample poses to a given model. Rather than relying solely on static heuristics or dataset-level statistics, our definition measures how well the model can approximate the desired tool invocation behavior before training, thereby capturing a performance-aware notion of difficulty.
@@
We aggregate all pairwise scores into a similarity matrix $S\in\mathbb{R}^{m\times m^{\prime}}$. To compute the best structural alignment between $A$ and $A^{\prime}$, we frame the task as a bipartite matching problem over the index sets $\mathcal{I}={1,\dots,m}$ and $\mathcal{J}={1,\dots,m^{\prime}}$, and apply the Hungarian algorithm [13] to identify the optimal matching $\mathcal{M}$ that maximizes the total alignment score:
@@
This score reflects the degree to which the model reproduces the reference tool invocation, normalized by the longer of the two sequences to account for missing or superfluous tool calls.
@@
Notably, there may be samples where $A_{1}=A$, meaning the refined answer is identical to the previous answer. We retain these samples to enable the model to learn that "no further changes are needed when the answer is perfect." This is a crucial aspect of establishing the model’s adaptive, iterative self-refinement, and we will provide further details in Section 3.3.
@@
3.3 Adaptive Self-Refine Iterative Inference
@@
We propose an adaptive self-refine procedure that enables ToolACE-R to iteratively refine its own output and autonomously determine when to halt the iteration. Details are shown in Figure 1 (right hand side) and Algorithm 1 in Appendix. Specifically, given a user query $q$ and candidate tools $T$, ToolACE-R first generates the tool invocation $A^{\prime}$. In each subsequent iteration, ToolACE-R leverages the answer from the previous iteration as context and refines it. As mentioned earlier, ToolACE-R has learned that when the answer is sufficiently accurate, it can produce the same answer in subsequent refinement. Therefore, the iteration continues until the answers from two consecutive iterations are identical. We refer to this procedure as adaptive self-refine, where the model itself determines when to stop the iterations. To prevent infinite iteration, we set a maximum iteration limit $n$. Through adaptive self-refine, ToolACE-R is able to improve its performance, particularly when facing difficult cases.
4 Experiments
4.1 Experimental Settings
@@
We leverage a subset of ToolACE [19] training data as our original training data. The dataset contains multiple types of tool calling data, including cases with multiple candidate tools and parallel function callings, in both Chinese and English. We select LLaMA3.1-8B-Instruct [2] for base model in our main experiments. We also conduct experiments on other base models, including Qwen2.5-Instruct-series [39] (0.5B, 1.5B, 3B and 7B) and Mistral-7B-Instruct-v0.2 [11], to validate the generalizability of ToolACE-R. We compare with the state-of-the-art API-based models, including GPT-series [^1] https://chatgpt.com and Gemini-2.5-Pro [^2] https://gemini.google.com/, as well as open-source models like DeepSeek-V3 [17], Llama3.1-8B/70B-Instruct [2] and Qwen2.5-7B-Instruct [39]. We also compare with fine-tuned tool calling models like ToolACE-8B [19], Hammer2.1-7B [16] and xLAM-2-8b-fc-r [41, 24].
@@
We conduct experiments on several representative benchmarks, including the Berkeley Function Call Leaderboard (BFCL) [38], ACEBench [5], API-Bank [15] and ToolAlpaca [33]. To minimize external interference and simplifying data preparation, we mainly evaluate performance on single-turn tool calling queries. For BFCL, we evaluate both the Non-live and Live subsets, which correspond to synthetic test cases and real-world scenarios, respectively. Each subset includes four categories: Simple, Multiple, Parallel, and Multiple Parallel. For ACEBench, we evaluate only the English normal category, excluding multi-turn cases. This includes the test samples in atom category and single-turn category. For API-Bank and ToolAlpaca, we focus on the first step tool calling samples, disregarding further tool or retrieval feedback. All performance is reported in terms of accuracy, where only cases with entirely correct tool calling are counted as correct.
Implementation Details.
Given resource constraints, we employ the parameter-efficient fine-tuning method LoRA [9]. All model modules are configured for LoRA fine-tuning, with a rank of 16 and an alpha value of 32. Training is performed with a global batch size of 64 and a learning rate of $1\times 10^{-4}$, following a cosine learning rate schedule with a warmup ratio of 0.1. For difficulty estimation, we set the generation temperature to 1.0 to encourage diversity and allow the model to explore a broader range of outputs. Model generates 8 times (i.e. $k=8$) for each sample to ensure robustness. During data selection, we set $\alpha=0$ and $\beta=0.9$, which means that we discard samples that are extremely easy (difficulty score equals to 0, i.e., the model always answer fully correctly) or too difficult (most attempts are total incorrect). For the constructed refinement data, we apply a loss mask to the first assistant turn of the conversation, as it may involve an incorrect tool invocation, making it unsuitable for the model to learn from this turn. During evaluation, we adopt greedy search to ensure stability. Unless otherwise specified, , the maximum iteration time for adaptive self-refine is set to $n=5$ .
4.2 Main Results
@@
Table 1 presents the comparison results on BFCL, including detailed results for each category and the overall average performance. The following key observations can be made:
  • The performance gap between API-based and open-source models is minimal. Large open-source models like DeepSeek-V3 and Llama3.1-70B-Inst show comparable or even superior overall performance to GPT models and Gemini-2.5-Pro.
  • Specialized fine-tuned models significantly benefit from domain-specific fine-tuning, allowing smaller models (7/8B) to compete with larger general models like Llama3.1-70B-Inst.
@@
To further assess the effectiveness of our adaptive self-refine method for iterative inference, we conducted experiments comparing with two other inference-time scaling approaches. The first, termed "Vanilla Self-Refine," always selects the last answer at each iteration as the final answer. The second method, "Self-Consistency at Itr n", chooses the majority answer from the answers collected during iterations 1 to n. The results of these inference methods on BFCL, are shown in Figure 2.
As observed, the performance of "Vanilla Self-Refine" fluctuates across iterations, indicating that the model does not always refine the answer correctly. In contrast, our "Adaptive Self-Refine" method demonstrates more stability across iterations, suggesting that adjacent identical answers serve as a reliable signal for stopping further iterations. Our statistics also show that the average iteration time of our method on BFCL is $2.4$ (when maximum iteration time $n=5$), indicating that most cases terminate early thus achieving consistent performance. Our method achieves competitive performance compared to the strong baseline "Self-Consistency at Itr 10", further validating advantages.
@@
Table 2: Accuracy comparison on ACEBench, API-Bank and ToolAlpaca benchmarks. “SR” is short for “Self-Refine”.
@@
Table 3: Ablation Study of our proposed modules on BFCL. “SR” is short for “Self-Refine”. The modules are progressively removed.
@@
4.3 Effects of Model-Aware Iterative Training
@@
In our main experiments, we set $\alpha=0$ and $\beta=0.9$ to discard samples that are either extremely simple or too difficult for model training. Here we experiment with more different choices of the two thresholds. We turn to another three kinds of choices, preserving simpler samples, preserving more difficult samples, or preserving samples in more medium difficulty. Table 5 displays the evaluation results on BFCL. As shown, setting $\alpha=-1$ and $\beta=0.9$, which includes even extremely simple cases (i.e., $\alpha=0$), leads to a slight performance drop while increasing training time due to the inclusion of a larger training corpus. In contrast, preserving only relatively difficult cases – by setting $\alpha=0.1$ and $\beta=2$ – results in a notable performance degradation across both categories, with an overall decrease of approximately 1 point. This suggests that samples beyond the model’s current capability may hinder effective learning and limit performance improvements based on its existing potential. We further evaluate a stricter setting ($\alpha=0.1,\beta=0.8$), which discards more relatively simple and difficult cases. This also results in performance drop, especially for Live category. The likely reason is that an excessive number of training samples are discarded (reducing the dataset to approximately half the size of our main setting) while retaining only a small portion of the harder examples.
@@
Table 4: Accuracy performance on BFCL when applying different threshods during model-aware data selection. We always select samples that their difficulty score satisfy $\alpha<D<\beta$.
@@
Table 5: Accuracy performance on BFCL when setting $k$ value in Eq. 5 as 88 and 11. “SR” is short for “Self-Refine”.
@@
Table 6: Accuracy performance on BFCL when using different selected data to train Llama3.1-8B-Inst, along with the data amount. “Qwen2.5-3/7B” is short for Qwen2.5-3/7B-Inst.
@@
Table 7: Accuracy performance of using different backbones on BFCL.
@@ 4.4 Further analysis
Figure 3: Accuracy performance and training amount of each iteration during training. Iteration 0 means the model before training.
Figure 4: Accuracy Performance of Qwen2.5-Series. “Base” refers to the Qwen2.5-Instruct models, while “+ ToolACE-R” uses our method.
@@
We further examine whether our method can be applied to models of varying sizes and whether performance improves as model size increases. For this, we use Qwen2.5-Instruct models of different sizes, including 0.5B, 1.5B, 3B, and 7B, as base models. The results, shown in Figure 4, display both the original performance and the performance after applying ToolACE-R. Applying ToolACE-R consistently improves performance across all sizes, particularly for smaller models like 0.5B and 1.5B. This demonstrates the effectiveness of ToolACE-R across a range of model sizes.
5 Conclusion

### v3

@@ Abstract
Tool learning, which allows Large Language Models (LLMs) to leverage external tools for solving complex user tasks, has emerged as a promising avenue for extending model capabilities. However, existing approaches primarily focus on data synthesis for fine-tuning LLMs to invoke tools effectively, largely ignoring how to fully stimulate the potential of the model. In this paper, we propose ToolACE-R, a novel framework that includes both model-aware iterative training and adaptive refinement for tool learning. ToolACE-R features a model-aware iterative training procedure that progressively adjust training samples based on the model’s evolving capabilities to maximize its potential. Additionally, it incorporates self-refinement training corpus which emphasizes LLM’s ability to iteratively refine their tool calls, optimizing performance without requiring external feedback. Furthermore, we introduce adaptive self-refinement for efficient test-time scaling, where the trained model can autonomously determine when to stop the process based on iterative self-refinement. We conduct extensive experiments across several benchmark datasets, showing that ToolACE-R achieves competitive performance compared to advanced LLMs. The performance can be further improved efficiently through adaptive self-refinement. These results highlight the effectiveness and generalizability of ToolACE-R, offering a promising direction for more efficient and scalable tool learning.
Introduction
@@
Tool learning, enabling Large Language Models (LLMs) to leverage external tools to address complex user requirements, has gained increasing attention. With tool integration, LLMs can access up-to-date information, perform intricate computations, and utilize third-party services, significantly expanding their capabilities beyond simple natural language communication with humans (Qu et al. 2024b). While tool invocation requires the LLMs to demonstrate strong understanding, reasoning, and instruction-following skills, customized fine-tuning is currently the dominant approach for enabling models to call external tools (Liu et al. 2024b, c; Patil et al. 2023; Qin et al. 2023).
@@
Due to the limited availability of high-quality data, existing research has primarily focused on developing effective and efficient methods for data synthesis using advanced models (Abdelaziz et al. 2024; Liu et al. 2024b, c; Wang et al. 2024b; Prabhakar et al. 2025). However, data synthesized by advanced models from different sources can lead to compatibility issues. Specifically, when synthesized samples exceed the model’s current knowledge, they may undermine the model’s performance or lead to hallucinations (Kang et al. 2024; Ren et al. 2024). As a result, determining which training data samples are appropriate for a given model remains an unresolved challenge.
@@
Additionally, the potential for maximizing a model’s intrinsic capabilities remains underexplored in the field of tool learning. On one hand, prior work has seldom investigated how to fully exploit existing training data through techniques such as data augmentation. On the other hand, scaling test-time computation – despite its demonstrated effectiveness in enhancing LLM reasoning (Brown et al. 2024; Snell et al. 2024) – has received limited attention in tool learning scenarios. A promising approach involves iteratively refining the model’s outputs and selecting the majority answer (Snell et al. 2024). However, user queries vary widely in complexity, ranging from simple to highly intricate. Scaling test-time computation indiscriminately is inefficient, as models can often correctly answer straightforward questions without the need for additional refinement. Therefore, an adaptive strategy for scaling test-time computation is needed to dynamically determine the appropriate level of computational effort required for each query.
@@
We have conducted extensive experiments on several representative tool-calling benchmarks, such as the Berkeley Function Call Leaderboard (BFCL) (Yan et al. 2024) and API-Bank (Li et al. 2023), to assess the effectiveness and efficiency of ToolACE-R. Experimental results demonstrate that ToolACE-R achieves competitive performance compared to advanced API-based models such as GPT-4o, and can be further enhanced through adaptive self-refinement. Additional analyses highlight the contributions of the proposed modules and examine the generalizability of our method across different model backbones and sizes.
The contributions of this work are summarized as follows:
@@
Tool learning methods generally fall into two categories, tuning-free and tuning-based approaches. Tuning-free methods employ prompting techniques without additional training (Mialon et al. 2023; Hsieh et al. 2023; Ruan et al. 2023). A notable example method is ReAct (Yao et al. 2023), which can alternate between reasoning and action for complex tasks. Recent work (Qu et al. 2024a) improves tool understanding by iteratively refining documentation. Tuning-based methods, which directly enhance tool calling ability via specialized fine-tuning, have recently gained much more attention (Qin et al. 2023; Schick et al. 2023; Patil et al. 2023; Tang et al. 2023; Abdelaziz et al. 2024). A large amount of them focus on data synthesis for improved generation procedures (Liu et al. 2024b, c; Wang et al. 2024b), often overlooking the model’s inherent potential. Zeng et al. (2025) also apply iterative training but based on DPO (Rafailov et al. 2023) which introduces more complexity. While some others integrate tool feedback for refinement (Du, Wei, and Zhang 2024; Wang et al. 2024a), our approach is more streamlined, directly refining outputs in an iterative manner without external or any textual feedback to maximize model efficiency.
@@
Selecting high-quality training samples is essential for fine-tuning LLMs (Albalak et al. 2024). A small set of high-quality data can effectively harness the model’s potential, rather than relying on large quantities of data (Zhou et al. 2023; Liu et al. 2023). While earlier works primarily emphasize general data quality aspects, such as diversity and complexity, recent studies advocate for model-specific data selection (Du, Zong, and Zhang 2023; Li et al. 2024). This kind of approach is based on the observation that data distributions that significantly deviate from the base model’s are challenging for the model to learn from and may even degrade the performance (Ren et al. 2024). Our method aligns with this perspective, defining a new metric to identify and select suitable training samples based on the model to be trained.
@@
Previous research has demonstrated that LLMs can refine their own generations through either self-feedback (Madaan et al. 2023; Weng et al. 2023) or external feedback (Qu et al. 2024c; Xu et al. 2024). Nevertheless, it remains challenging for LLMs to assess the correctness of their refined output autonomously (Huang et al. 2024), where scaling test-time compute with iterative refinement still often depends on post-processing techniques, such as majority voting (Snell et al. 2024). In this work, we propose a simple yet effective method that enables LLMs to adaptively self-refine their outputs, improving scaling efficiency.
@@
Methodology
Problem Formulation
Given a user query $q$ and a set of candidate tools $T=\{t_{1},t_{2},\dots,t_{n}\}$, the objective of tool learning is to generate the correct tool invocations. This includes selecting the most appropriate tools and extracting the suitable parameter values. Specifically, the goal is to produce:
@@
ToolACE-R is trained with a model-aware iterative training pipeline and can iteratively self-refine during inference in an adaptive manner. Figure 1 shows the overall procedure.
Model-Aware Iterative Training
@@
We introduce a model-aware difficulty metric that quantifies the inherent challenge a specific training sample poses to a given model. Rather than relying solely on static heuristics or dataset-level statistics, our definition measures how well the model can approximate the desired tool invocation behavior before training, thereby capturing a performance-aware notion of difficulty.
@@
We aggregate all pairwise scores into a similarity matrix $S\in\mathbb{R}^{m\times m^{\prime}}$. To compute the best structural alignment between $A$ and $A^{\prime}$, we frame the task as a bipartite matching problem over the index sets $\mathcal{I}={1,\dots,m}$ and $\mathcal{J}={1,\dots,m^{\prime}}$, and apply the Hungarian algorithm (Kuhn 1955) to identify the optimal matching $\mathcal{M}$ that maximizes the total alignment score:
@@
This score reflects the degree to which the model reproduces the reference tool call, normalized by the longer of the two sequences to account for missing or superfluous tool calls.
@@
Notably, there may be samples where $A_{1}=A$, meaning the refined answer is identical to the previous answer. We retain these samples to enable the model to learn that ”no further changes are needed when the answer is perfect.” This is a crucial aspect of establishing the model’s adaptive, iterative self-refinement, and we will provide further details in later section.
@@
Adaptive Self-Refine Iterative Inference
@@
We propose an adaptive self-refine procedure that enables ToolACE-R to iteratively refine its own output and autonomously determine when to halt the iteration. Details are shown in Figure 1 (right hand side). Specifically, given a user query $q$ and candidate tools $T$, ToolACE-R first generates the tool invocation $A^{\prime}$. In each subsequent iteration, ToolACE-R leverages the answer from the previous iteration as context and refines it. As mentioned earlier, ToolACE-R has learned that when the answer is sufficiently accurate, it can produce the same answer in subsequent refinement. Therefore, the iteration continues until the answers from two consecutive iterations are identical. We refer to this procedure as adaptive self-refinement, where the model itself determines when to stop the iterations. To prevent infinite iteration, we set a maximum iteration limit $n$. Through adaptive self-refinement, ToolACE-R is able to improve its performance, particularly when facing difficult cases.
Experiments
Experimental Settings
@@
We use a subset of ToolACE (Liu et al. 2024b) training data as our training data, where we only retain the samples with single-turn tool calling for simplicity. We select LLaMA3.1-8B-Instruct (AI@Meta 2024) as base model in our main experiments. We also conduct experiments on Qwen2.5-Instruct-series (Yang et al. 2024) (0.5B, 1.5B, 3B and 7B) and Mistral-7B-Instruct-v0.2 (Jiang et al. 2023), to validate the generalizability of ToolACE-R. We compare with the state-of-the-art API-based models, including GPT-series and Gemini-2.5-Pro, as well as open-source models like DeepSeek-V3 (Liu et al. 2024a), Llama3.1-8B/70B-Instruct (AI@Meta 2024) and Qwen2.5-7B-Instruct (Yang et al. 2024). We also compare with fine-tuned tool calling models like ToolACE-8B (Liu et al. 2024b), Hammer2.1-7B (Lin et al. 2024) and xLAM-2-8b-fc-r (Zhang et al. 2024; Prabhakar et al. 2025).
@@
We conduct experiments on several representative benchmarks, including the Berkeley Function Call Leaderboard (BFCL) (Yan et al. 2024), ACEBench (Chen et al. 2025), API-Bank (Li et al. 2023) and ToolAlpaca (Tang et al. 2023). To minimize external interference and simplifying data preparation, we mainly evaluate performance on single-turn tool calling queries. For BFCL, we evaluate the Non-live and Live subsets, which correspond to synthetic test cases and real-world scenarios, respectively. Each subset includes four categories: Simple, Multiple, Parallel, and Multiple Parallel. For ACEBench, we evaluate only the English normal category, excluding multi-turn cases. This includes the test samples in atom category and single-turn category. For API-Bank and ToolAlpaca, we focus on the first step tool calling samples, disregarding further tool or retrieval feedback. All performance is reported in terms of accuracy, where only cases with entirely correct tool calling are counted as correct.
@@
Implementation Details.
We employ parameter-efficient fine-tuning method LoRA (Hu et al. 2022) given resource constraints. All model modules are configured for LoRA fine-tuning, with a rank of 16 and an alpha value of 32. Training is performed with a global batch size of 64 and a learning rate of $1\times 10^{-4}$, following a cosine learning rate schedule with a warmup ratio of 0.1. For difficulty estimation, we set the generation temperature to 1.0 to encourage diversity and allow the model to explore a broader range of outputs. Model generates 8 times (i.e. $k=8$) for each sample to ensure robustness. During data selection, we set $\alpha=0$ and $\beta=0.9$, which means that we discard samples that are extremely easy (difficulty score equals to 0, i.e., the model always answer fully correctly) or too difficult (most attempts are total incorrect). For the constructed refinement data, we apply a loss mask to the first assistant turn of the conversation, as it may involve an incorrect tool invocation, making it unsuitable for the model to learn from this turn. During evaluation, we adopt greedy search to ensure stability. Unless otherwise specified, , the maximum iteration time for adaptive self-refinement is set to $n=5$ .
Main Results
@@
Table 1 presents the comparison results on BFCL, including detailed results for each category. The following key observations can be made:
  • The performance gap between API-based and open-source models is minimal. Large open-source models like DeepSeek-V3 show comparable or even superior overall performance to GPT models and Gemini-2.5-Pro.
  • Specialized fine-tuned models significantly benefit from domain-specific fine-tuning, where smaller models (7/8B) are allowed to compete with larger general models like Llama3.1-70B-Inst.
@@
To further assess the effectiveness of our adaptive self-refine method for iterative inference, we conducted experiments comparing with two other inference-time scaling approaches. The first, termed ”Vanilla Self-Refine,” always selects the last answer at each iteration as the final answer. The second method, ”Self-Consistency at Itr n”, chooses the majority answer from the answers collected during iterations 1 to n. The results of these inference methods on BFCL, are shown in Figure 2.
As observed, the performance of ”Vanilla Self-Refine” fluctuates across iterations, indicating that the model does not always refine the answer correctly. In contrast, our ”Adaptive Self-Refine” method demonstrates more stability across iterations, suggesting that adjacent identical answers serve as a reliable signal for stopping further iterations. Our statistics also show that the average iteration time of our method on BFCL is $2.4$ (when maximum iteration time $n=5$), indicating that most cases terminate early thus achieving consistent performance. Our method achieves competitive performance compared to the strong baseline ”Self-Consistency at Itr 10”, further validating advantages.
@@
Table 2: Accuracy comparison on ACEBench, API-Bank and ToolAlpaca. “SR” is short for “Self-Refine”.
@@
Table 3: Ablation Study of our proposed modules on BFCL. “SR” is short for “Self-Refine”.
@@
Effects of Model-Aware Iterative Training
@@
In our main experiments, we set $\alpha=0$ and $\beta=0.9$ to discard samples that are either extremely simple or too difficult for model training. Here we experiment with more different choices of the two thresholds. We turn to another three kinds of choices, preserving simpler samples, preserving more difficult samples, or preserving samples in more medium difficulty. Table 4 displays the evaluation results on BFCL. As shown, setting $\alpha=-1$ and $\beta=0.9$, which includes even extremely simple cases (i.e., $\alpha=0$), leads to a slight performance drop while increasing training time due to the inclusion of a larger training corpus. In contrast, preserving only relatively difficult cases (by setting $\alpha=0.1$ and $\beta=2$) results in a notable performance degradation across both categories, with an overall decrease of approximately 1%. This suggests that samples beyond the model’s current capability may hinder effective learning and limit performance improvements based on its existing potential. We further evaluate a stricter setting ($\alpha=0.1,\beta=0.8$), which discards more simple and difficult cases. This also results in performance drop, especially for Live category. The likely reason is that an excessive number of training samples are discarded (reducing the dataset to approximately half the size of our main setting) while retaining only a small portion of the harder examples.
@@
Table 4: Performance on BFCL when applying different threshods during model-aware data selection. We always select samples that their difficulty score satisfy $\alpha<D<\beta$.
@@
Table 5: Performance on BFCL when setting $k$ value in Eq. 5 as 88 and 11. “SR” is short for “Self-Refine”.
@@
Table 6: Performance on BFCL when using different selected data to train Llama3.1-8B-Inst, along with the data amount. “Qwen2.5-3/7B” is short for Qwen2.5-3/7B-Inst.
@@
Table 7: Performance of using different models on BFCL.
@@
Figure 3: Performance and training amount of each iteration during training. Iteration 0 means the model before training.
Figure 4: Performance of Qwen2.5-Series. “Base” refers to the Qwen2.5-Instuct models without any training, while “+ ToolACE-R” uses our method.
@@
We further examine whether our method can be applied to models of varying sizes and whether performance improves as model size increases. We use Qwen2.5-Instruct models of different sizes, including 0.5B, 1.5B, 3B, and 7B, as base models. The results, shown in Figure 4, display both the original performance and the performance after applying ToolACE-R. Applying ToolACE-R consistently improves performance across all sizes, particularly for smaller models like 0.5B and 1.5B. This demonstrates the effectiveness of ToolACE-R across a range of model sizes.
Conclusion